OpenStudy (anonymous):

Which expressions would complete this equation so that the equation has no solution? 6(4x – 3) – 10x = Choose exactly two answers that are correct. A. 14x – 15 B. 2(7x – 10) C. 15x – 14 D. 2(7x – 9)

OpenStudy (anonymous):

A & B - A & B have the same slope = 14 ==> They run parallel to each other, and to f(x) - D(x) have the same slope m =14 too but f(x) and function D are identical functions. So with every x, we have f(x) = D(x)